The equity shares of a firm in the current stock market has been traded at Rs60 per share. The price earnings ratio is 10 times. The Dividend payout ratio is 75%.

The total number of shares issued and outstanding as on date are 100000 equity shares of Rs10 each. Book value of each share is Rs40

Describe and Compute - Earnings per share, return on equity


1
Expert's answer
2021-10-02T10:48:56-0400

If the price earnings ratio is 10 times, then the earnings per share is 60/10 = Rs 6.

The dividend per share is 6×0.75 = Rs 4.5.

Total net income is: 6×100,000 = Rs 600,000.

Shareholders equity is: 40×100,000 = Rs 4,000,000.

ROE = 600,000/4,000,000 = 0.15.